OpenStudy (anonymous):

does the sqr root sign covering a whole term of a fraction in which both the numerator and denominator are themselves squared mean I can rewrite it by simply cancelling out the sqr root and squared signs ? wait a minute Im writing the expression :)

OpenStudy (anonymous):

\[\sqrt{2^2/4^2} = 1/2 ?\]

OpenStudy (***[isuru]***):

yep.!!!

OpenStudy (anonymous):

\[\sqrt{\frac{a}{b}} = \frac{\sqrt{a}}{\sqrt{b}}\] I think that's what you're asking.

OpenStudy (anonymous):

yes both !!

OpenStudy (anonymous):

Yes, you can do that.

OpenStudy (anonymous):

well its being difficult because I can also get that sum to equal \[\sqrt{1/4}\]

OpenStudy (anonymous):

and can you see where im getting mixed up ?

OpenStudy (***[isuru]***):

\[\sqrt{\frac{ 2^{2} }{4^{2}}} = \frac{\sqrt{2^{2}} }{ \sqrt{4^{2}} } = \frac{ 2^{2 \times \frac{ 1 }{ 2 } }}{ 4 ^{2 \times \frac{ 1 }{ 2 }} }\]

OpenStudy (anonymous):

ok....

OpenStudy (***[isuru]***):